OpenStudy (anonymous):

PLEASE HELP! What is the y coordinate of the y−intercept of the line that passes through the points (−4, −1) and (4, 5)?

OpenStudy (anonymous):

Okay first you need to calculate the equation of the line, for that you need the gradient of the line passing through the two points. \[m = (y _{2} - y_{1}) \div (x _{2} - x_{1})\] where m is the gradient. Using this, you can select any of the points and compute gradient. The gradient is, m = 3/4 Now you use the general form of a linear equation to find the equation of the line: \[y - y_{1} = m(x - x_{1})\] You leave x and y as they are and select ONE of the point sets to be x1 and y1 Hence you get y-5 = 0.75(x-4) y-5 = 0.75x - 3 y = 0.75x + 2 Remember, on the y axis, x=0 So the y-intercept is 2

OpenStudy (anonymous):

IM SO GRATEFUL ! THANK YOU SO MUCH! :D

OpenStudy (anonymous):

no problem :)